two inductors with both the self-inductance and the induced emf given for each. Which inductor has the greater current flowing through it?

2 H 2V 1H 4 V Inductor A Inductor B

A) inductor A

B) inductor B

C) equal

D) there is insufficient information
Answer #1

EMF = L di/dt

di/dt = V/L

for inductor A

di/dt = 2/2 = 1 A/s

FOR INDUCTOR B

di/dt = 4/1 = 4 A/s

so inductor B has greater current flowing through it

so option B) is the correct answer